peter is 36 years old. Peter is twice as old as jun was when peter was as old as jun is now. How old is Jun?
Answer by ikleyn(52786)

.
Peter is 36 years old. Peter is twice as old as Jun was when Peter was as old as Jun is now. How old is Jun?
~~~~~~~~~~~~~~~~~~~~~~

Let J be the Jun's age now.


How many years ago Peter was as old as Jun is now ? --- ANSWER :  36-J years ago.


How old was Jun then ? --- ANSWER :  J - (36-J)  = 2J-36  years old.


Now we can write the setup equation, as we read the problem


    36 = 2*(2J-36).


Simplify and find J from this equation


    36 = 4J - 72

    36 + 72 = 4J

      108   = 4J

       J    = 108/4 = 27.


ANSWER.  Jun is 27 years old.


CHECK.   Peter was as old as Jun is now 36-27 = 9 years ago.

         9 years ago Jun was 27-9 = 18 years.

         Peter's age now (36) is twice Jun's age then  (9 years ago Jun was 18 years old.

                            ! Correct !

Solved.

-----------------

It is a twisted age word problem.

To see many other similar problems solved, look at the lesson
    - Really intricate age word problems
in this site.


Also,  you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lesson is the part of this online textbook under the topic "Age word problems".


Save the link to this online textbook together with its description

Free of charge online textbook in ALGEBRA-I
https://www.algebra.com/algebra/homework/quadratic/lessons/ALGEBRA-I-YOUR-ONLINE-TEXTBOOK.lesson

to your archive and use it when it is needed.